The Scampini Supplies Company recently purchased a new delivery truck. The new truck costs $25,000, and it is expected to generate after-tax cash flows, including depreciation, of $6,500 per year. The truck has a 5-year expected life. The expected year-end abandonment values (salvage values after tax adjustments) for the truck are given below. The company's WACC is 11%.

Year Annual After-Tax Cash Flow Abandonment Value
0 ($25,000) -
1 6,500 $20,000
2 6,500 15,500
3 6,500 13,500
4 6,500 7,500
5 6,500 0

What is the truck's optimal economic life? Round your answer to the nearest whole number.

   year(s)
Answer #1
Present Value (PV) of Cash Flow:
(Cash Flow)/((1+i)^N)
i=Discount Rate=WACC=11%=0.11
N=Year of Cash Flow
N A B=A/(1.11^N) C=B-25000 D E=D/(1.11^N) F G=C+F H(Using PMT function with Rate=11%,Nper=N,PV=-G)
End of year Abandonment Value PV of abandonment Value PV of net capital cost Annual Cash Flow PV of annual Cash flow Cumulative PV of annual Cash flow PV of Net Cash Flow Equivalent annual Cashflow
0
1 $20,000 $18,018 ($6,982) $6,500 $             5,856 $    5,856 ($1,126) ($1,250)
2 $15,500 $12,580 ($12,420) $6,500 $             5,276 $ 11,131 ($1,288) ($752)
3 $13,500 $9,871 ($15,129) $6,500 $             4,753 $ 15,884 $755 $309
4 $7,500 $4,940 ($20,060) $6,500 $             4,282 $ 20,166 $106 $34
5 $0 $0 ($25,000) $6,500 $             3,857 $ 24,023 ($977) ($264)
PV of net Cash flow is highest in Year3
Equivalent annual cash flow is highest in year 3
OPTIMAL ECONOMIC LIFE 3 Years